MONT 104N – Modeling the Environment First Paper Assignment In Two Stages; September 3, 2019

Sources

For this assignment, I would like you to begin by looking up and reading two reviews of our cluster common text The Uninhabitable Earth. Both of these are easily findable through the HC Library site, and also directly by Google. (See me or Ms. Merolli, our Science Librarian, if you have any trouble tracking them down.) The first is titled “The Uninhabitable Earth review: Astonishingly reductive view of impending disaster,” written by John Gibbons and published in The Irish Times on Feb. 23, 2019. Here’s a portion so you get the idea:

“What I found disappointing is that his grim journey of discovery does not yet appear to have widened his circle of empathy beyond the fate of humanity. Yet already, the unmistakable scent of death is all around us. Scientists estimate that between 150-200 species of plant, insect, bird and mammal become extinct every 24 hours as ecosystems unravel.

Yet the destruction of the natural world and collapse in global biodiversity only appears to concern Wallace-Wells insofar as it impacts on one species. One of his all-too-rare forays into looking at climate impacts on non-humans involves the shocking mass death of hundreds of thousands of an Asian dwarf antelope, the saiga, in a single incident in May 2015. Elevated temperatures had caused a usually benign bacteria to suddenly turn deadly. Once again, the author only filters this as a potentially cautionary tale for humans.”

The second is the review of The Uninhabitable Earth written by John Lanchester, published in The New York Times, April 12, 2019. (This was paired with another book on a related topic, and the review actually covers both books. You only need to look at the part about our common text, though you may find the rest interesting too.) Again here’s a sample:

“Pessimism [in the face of climate change] would be an ethical catastrophe. It leads only to despair, despair to inaction, and inaction to a future world David Attenborough has described as “the collapse of our civilizations and the extinction of much of the natural world.” To avoid the most terrible possible versions of our future, we have to stay positive; it’s the only moral response to this crisis. And there are grounds to do so, as David Wallace-Wells argues in his brilliant new book, “The Uninhabitable Earth: Life After Warming”: “We have all the tools we need, today, to stop it all: a carbon tax and the political apparatus to aggressively phase out dirty energy, a new approach to agricultural practices and a shift away from beef and dairy in the global diet; and public investment in green energy and carbon capture.”

Global emissions could be cut by a third if the richest 10 percent of humanity cut their use of energy to the same level as affluent, comfortable Europe. One prospective technique to scrub carbon from the atmosphere would cost $3 trillion a year, a colossal amount but significantly less than the current level of subsidies paid out globally for fossil fuel, estimated at $5 trillion. Taken all in all, solutions are “obvious” and “available.” The only obstacle to implementing them is political will.”

The Assignment

This assignment will have two parts or stages, along the lines of the “They Say, I Say” paradigm we will discuss in class:

• Stage 1–The “They Say” part: You will write about two double-spaced pages to paraphrase and summarize the opinions presented the two reviews identified above. Please don’t add your own opinion or express a preference for either reviewer’s point of view yet. This first part of the assignment is geared toward seeing how well and carefully you read these reviewers’ words and how faithfully you can present the ideas in these two sources in your own words.

• The first stage will be due by email to jlittle@holycross.edu no later than 5:00pm on Monday, September 16.

• I will return the Stage 1 later that week, by Friday September 20 with comments.

• Stage 2–The “I Say” part: Complete an essay of three to five double-spaced pages in all, using your Stage 1, revised as needed, followed by your own take on these points about the book raised by the reviewers. Be honest – you can prefer one or the other, you might think both reviews missed the boat in some way, or you might even agree with both of them. Here the goal will be to come up with a judgment of your own, supported as appropriate by quotations from the book, the reviews, etc. Reasonable people can come to completely different conclusions, and what this part of the assignment is about is getting you to think through a position of your own and articulate it convincingly.

• The second stage will be due, by email again to jlittle@holycross.edu, no later than 5:00pm on Wednesday, September 25.

Guidelines for this assignment

Follow these specifications:

1. Stage 1 is all about conveying what other people have said about the book. Again, don’t prematurely insert your own opinions and try not to let whatever opinions you have about these issues lead you into misrepresentations. Your presentation of the points of view presented in the two reviews should be clear, complete, and it should correctly convey what those authors were trying to say.

2. If you use direct quotations from either review, identify which one you are drawing from with a parenthetical reference in the form (author’s name). If you quote from Wallace-Wells, use his name, and give the page number where you found the quotation.

3. You should not need to consult any other sources beyond the two reviews above and the book. If you really think you need to, please consult with Prof. Little and get approval.

4. When you get to Stage 2, the paper as a whole should be well-organized and “flow” from one point to the next. Do not repeat points unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

5. Your sentences should be clear, readable, and not overly wordy. You should aim fora lively style and the writing should be in your own voice. But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

6. Errors in mechanics (spelling, punctuation, grammar) and usage (word choice) should not annoy the reader (i.e. Prof. Little). (A word of warning: Your reader can be easily annoyed by this sort of thing at times!) Try as much as possible to avoid those mistakes.